OpenStudy (anonymous):

write the equation of a line perpendicular to m= -1/3, (-2,-5)

OpenStudy (texaschic101):

A perpendicular line is going to have the negative reciprocal slope. So, the negative reciprocal slope of -1/3 is 3/1 or just 3. y - y1 = m(x - x1) using slope 3 and points (-2,-5) y - (-5) = 3(x - (-2) y + 5 = 3(x + 2) y + 5 = 3x + 6 y = 3x + 6 - 5 y = 3x + 1

OpenStudy (anonymous):

ok cool that helped a lot!! :) THANKS

OpenStudy (texaschic101):

anytime :)